Hi there, Im looking to find the sum of the first four terms of this sequence
-18,9,-9/2

Answer by macston(5194)

Multiply each term by -1/2:
36(-1/2)=-18
-18(-1/2)=9
9(-1/2)=-9/2
-9/2(-1/2)=9/4
9/4(-1/2)=-9/8
...